### v1

# Remote Sensing Reveals Adoption of Sustainable Rice Farming Practices Across Punjab, India

Abstract
Rice cultivation consumes 24-30% of global freshwater, creating critical water management challenges in major rice-producing regions. Sustainable irrigation practices like direct seeded rice (DSR) and alternate wetting and drying (AWD) can reduce water use by 20-40% while maintaining yields, helping secure long-term agricultural productivity as water scarcity intensifies—a key component of the Zero Hunger Sustainable Development Goal. However, limited data on adoption rates of these practices prevents evidence-based policymaking and targeted resource allocation. We developed a novel remote sensing framework to monitor sustainable water management practices at scale in Punjab, India—a region facing severe groundwater depletion of 41.6 cm/year. To collect essential ground truth data, we partnered with the Nature Conservancy’s Promoting Regenerative and No-burn Agriculture (PRANA), which trained approximately 1,400 farmers on water-saving techniques while documenting their field-level practices. Using this data, we created a classification system with Sentinel-1 satellite imagery that separates water management along sowing and irrigation dimensions. Our approach achieved a 78% F1-score in distinguishing DSR from traditional puddled transplanted rice without requiring prior knowledge of planting dates. We demonstrated scalability by mapping DSR adoption across approximately 3 million agricultural plots in Punjab, with district-level predictions showing strong correlation (Pearson=0.77, RBO= 0.77) with government records. This study provides policymakers with a powerful tool to track sustainable water management adoption, target interventions, and measure program impacts at scale.
Index Terms—  Deep Learning, Remote Sensing, GeoAI, Rice Agriculture, Applied Machine Learning, Sustainable Agriculture, Water Management
1 Introduction
Rice is a staple food for half of the world’s population, but traditional paddy cultivation requires enormous quantities of water—2,000–5,000 liters per kilogram of rice produced [1]. As climate change intensifies water scarcity in many regions, sustainable water management in rice cultivation becomes essential for ensuring long-term food security, critical to addressing the U.N. Zero Hunger Sustainable Development Goal (SDG). Two water-saving practices have emerged as particularly promising: Direct Seeded Rice (DSR) and Alternate Wetting and Drying (AWD). DSR eliminates the need to raise and transplant seedlings by directly sowing seeds into the field, reducing water consumption by 20-40% [2, 1], relative to traditionally practiced puddled transplanted rice (PTR). AWD involves periodically allowing fields to dry before re-flooding, reducing water use by up to 30% compared to traditional continuous flooding (CF) [3]. When properly managed, these techniques maintain or even enhance yields, presenting a balanced approach for sustainable rice production [4].
Despite the promise of these practices, their scale of adoption remains poorly understood. This knowledge gap severely hampers evidence-based policymaking, resource allocation, and intervention planning for several reasons:
  • Ground surveys are labor-intensive, expensive, and often constrained by limited access to remote fields
  • Manual methods struggle to differentiate between traditional and newly adopted practices
  • Adoption patterns are not routinely tracked in national or sub-national surveys
  • Understanding current adoption helps target resources where they’re most needed
Fig. 1: Mapping DSR predictions across Punjab for 2024: A) Map of Punjab showing DSR adoption distributions with district boundaries. B) Comparison between model predictions and government records of DSR adoption by district, ordered by acres of DSR activity, as estimated by the Punjab government. C) Correlation analysis (Pearson=0.79) between model predictions and government records with outlier and important districts labeled.
Remote sensing offers a potential solution for monitoring water management practices at scale. While satellite-based methods have successfully mapped rice cultivation using multispectral optical and synthetic aperture radar (SAR) approaches [5, 6, 7], detecting specific water management practices remains challenging. Existing studies rely heavily on prior knowledge of planting dates [8, 9] or use coarse resolution data [10], limiting their scalability in regions where planting dates vary significantly and in regions dominated by smallholder farming.
We focused our analysis on Punjab, India, a major rice-producing region facing severe water stress. Punjab contributes 10.42% of India’s rice production [11], but faces groundwater depletion of 41.6 cm/year [12]. This crisis stems from the region’s semi-arid climate, rice-wheat cropping systems that increased irrigation demand, and groundwater extraction facilitated by free electricity [13, 14]. Approximately 73% of irrigation water comes from rapidly depleting groundwater sources [15]. To address these challenges, we partnered with The Nature Conservancy’s Promoting Regenerative and No-burn Agriculture (PRANA) project [^1] https://www.nature.org/en-us/about-us/where-we-work/india/our-priorities/prana/. PRANA trained approximately 150,000 farmers across Punjab on sustainable water management methods, while collecting detailed field-level data from about 1,400 participants. This included documentation of sowing methods, irrigation practices, and precise field boundaries–critical ground truth data for our remote sensing approach. Our study introduces a dimensional framework for monitoring water management practices using Sentinel-1 satellite imagery. By separating practices along sowing (DSR vs. PTR) and irrigation (AWD vs. CF) dimensions, we leverage their distinct temporal signatures while avoiding the limitations of simultaneous detection. Importantly, our approach does not require prior knowledge of cropping calendars, enhancing its scalability for large-scale monitoring.
Fig. 2: Water-saving Interventions: A) AWD measuring pipe being installed in a demonstration field. B) AWD pipe after installation. C) Photograph of a sample field practicing DSR taken during the tillering stage; note the lack of flooding in the field.
Our main contributions include:
  • A dimensional classification framework that separates water management practices along sowing and irrigation dimensions
  • A robust model using Sentinel-1 time-series imagery that achieves 78% F1-score in classifying DSR despite significant variability in planting dates
  • A scalable plot-level method demonstrated across approximately 3 million agricultural plots in Punjab, with aggregated district-level predictions correlating strongly with government records
2 Background
2.1 Study Area
Our study focuses on Punjab, India (Fig. 5), a semi-arid region where rice dominates agriculture, covering 87% (3.13 million hectares) of farmland during the 2022-23 monsoon season [16]. In 2023, Punjab produced 14.36 million tons of rice—10.4% of India’s production and approximately 3.3% of global output [11]. The region typically employs a wheat-rice double-cropping system using PTR and CF methods[17].
This agricultural intensity has created a severe water crisis, with 73% of irrigation water drawn from rapidly depleting underground aquifers. The 2022 Central Ground Water Board report classified 114 of Punjab’s 150 blocks as over-exploited, with an additional 19 blocks categorized as critical or semi-critical [18]. Although the state government has implemented conservation measures, including legislation restricting early rice transplanting [15], groundwater levels continue to decline at an alarming rate of 41.6 cm/year, expected to drop below 300 metres by 2039 [12, 18].
Fig. 3: Temporal view of water management practices. Y-axis represents water level in a rice paddy, and x-axis shows growth stages. The growing season naturally separates into two phases: initial (left of dotted line) and main growth (right). Blue arrows indicate irrigation events, while yellow arrows denote sowing. The different practices create distinct temporal signatures, with the vertical dotted line marking the boundary between sowing and irrigation dimensions.
Fig. 4: Dimensional divisions of practices. A) Plots divided by sowing and irrigation dimensions, with number of plots in parentheses. B) Classification of irrigation practices (grouping Control and DSR as CF). C) Classification of sowing practices (grouping Control and AWD as PTR). Dotted red lines in B) and C) show classes are separated in a task-dependent manner.
2.2 Challenges
While satellite-based monitoring offers a potential solution, several technical challenges have limited its application:
2.2.1 Water Management Practices and Separability
Water-saving practices in rice cultivation occur at different crop stages and create distinct water level patterns. DSR affects water management during early sowing and tillering phases, while AWD impacts water dynamics from tillering through harvest.
As shown in Figure 4, these practices create unique temporal signatures in water levels and soil moisture that can be detected via Synthetic Aperture Radar (SAR) remote sensing [19, 8, 9, 10]. PTR maintains consistently high water levels, while DSR shows a sharp initial decline after pre-sowing irrigation. Similarly, CF maintains flooded fields throughout the season, whereas AWD creates cyclical wet-dry patterns.
2.2.2 Dimensional Classification Approach
Attempting to classify all water management combinations simultaneously poses substantial challenges due to overlapping characteristics. We address this by separating practices into two independent dimensions: sowing dimension, distinguishing between DSR and PTR practices and irrigation dimension, separating AWD from CF practices.
This dimensional approach, illustrated in Figure 4, allows us to cleanly separate practices and collect training data that isolates specific dimensions of interest. Control plots employ PTR and CF practices, DSR plots practice DSR and CF, and AWD plots practice PTR and AWD. Plots practicing mixed combinations were excluded to avoid confounding effects.
2.2.3 Planting Date Heterogeneity
An additional challenge for satellite-based monitoring is the significant variation in planting schedules. Previous studies have relied on synchronizing satellite data with known planting dates [8, 9] – a luxury unavailable when monitoring millions of plots. In our study area, planting dates span up to 110 days, with the majority of DSR plots typically sown first, followed by control plots and finally AWD plots (Fig. S3). This temporal variability introduces natural differences between classes that must be accounted for, and in fact is exploited, in our modeling approach.
2.2.4 Plot-level Predictions
To be able to provide fine-grained insights, we must be able to both identify individual plot boundaries as well as make predictions about these monolithic entities. We employ Fields of the World (FTW) algorithm [20] to automate the extraction of plot boundaries, enabling us to make predictions at an individual plot level.
Our research question thus becomes: How effectively can we classify water management practices by separating them along sowing and irrigation dimensions using satellite imagery, despite significant heterogeneity in planting dates, while predicting at the plot-level?
3 Materials and Methods
3.1 Ground Truth Data Collection
To obtain reliable training data, we partnered with The Nature Conservancy’s PRANA project, which trains farmers on sustainable water management practices across Punjab. Between May and December 2024, local field teams collected detailed plot-level data from PRANA intervention sites across 18 districts. The resulting dataset includes 452 AWD plots, 420 DSR plots and 411 PTR plots employing CF. Their average sizes are 4.2, 2.7 and 2.1 acres, respectively. Refer to Fig. S1 for details of AWD measuring devices and visuals of DSR plotsin the PRANA program.
For each plot, we documented sowing dates, transplantation dates (for PTR/AWD), irrigation schedules, crop varieties, harvesting dates, and precise field boundaries. Field delineations were conducted through on-site visits using custom software and Google Earth Pro. We selected only plots with clearly visible boundaries and a minimum area of 2000 square meters to ensure sufficient quality in satellite data analysis. A 10-meter negative buffer was applied to all boundaries to eliminate edge effects while preserving adequate pixels for analysis.
3.2 Satellite Data and Processing
We leveraged Sentinel-1 Synthetic Aperture Radar (SAR) data to monitor water management practices across the growing season (April-December 2024). We acquired 108 images from ascending orbit — 17 $\times$ 3 from row 27222An on-board anomaly on Sentinel-1A on June 27-28 led to irrecoverable loss of data and therefore a missing timestep in the row 27 and 19 $\times$ 3 from row 100 — covering the entire state of Punjab. Through experimentation, the ascending orbit dataset was shown to have higher performance and was selected for training and inference, focusing on the time period May 1 - December 15, the maximum duration of the kharif growing season for rice in Punjab. For details please refer to section 4.2.
Fig. 5: Map of Punjab showing the study area, Sentinel-1 satellite coverage (ascending and descending orbits), and locations of surveyed plots. Insets show sample plots at different zoom levels, with the innermost inset displaying delineated fields from three classes: Control (plots that use puddled transplanting and continuous flooding), alternate wetting and drying (AWD), and direct seeded rice (DSR).
SAR data was specifically selected for its cloud-penetrating capabilities, a crucial advantage during the monsoon season when optical imagery is frequently compromised. Previous have shown SAR’s effectiveness in capturing diverse rice agriculture metrics, including cultural type classification, phenological parameter detection, growing season estimation [21, 22, 23], and differentiation between multiple cropping cycles and field-level tillage practices [7, 24]. To account for environmental factors influencing irrigation decisions, we integrated ERA5-Land climate reanalysis data [25], incorporating daily temperature and precipitation values as contextual variables in our analysis framework.
All SAR imagery was processed using the SNAP toolbox [26] to extract time series information at 10-meter spatial resolution. We extract Sentinel-1 VV and VH polarization bands from multiple orbit passes over Punjab. For each image, we apply standard preprocessing steps: orbit file application, noise removal, radiometric calibration, speckle filtering using multi-temporal refined-Lee filter [27], and terrain correction to derive terrain-geocoded backscatter coefficient ($\gamma_{0}$) at 10-meter resolution [26].
For each agricultural plot polygon in the training corpus, we calculate mean backscatter values for every timestep, generating a timeseries per plot. This process serves both as speckle noise reduction for the SAR imagery, and as a summary of plot-level conditions. To account for temporal gaps between satellite passes and enable consistent analysis, we apply temporal smoothing with a 1-D Gaussian filter (standard deviation = 0.5) followed by fitting a third-degree polynomial spline. This allows us to sample the time series at regular intervals despite variable satellite acquisition dates–steps that are crucial in dealing with missing data in the timeseries. We additionally create a VV/VH band ratio time series, which has proven effective for rice mapping in previous studies [9, 28, 29].
3.3 Feature Engineering
We develop two complementary sets of features:
Hand-crafted Features: For each band (VV, VH, VV/VH), we extract temporal characteristics that include the position and amplitude of the first $k$ troughs, crests, and inflection points; the total number of troughs and crests; summary statistics (mean, minimum, maximum amplitude); parameters from Gaussian kernel fitting to the VV/VH ratio (amplitude, peak day, standard deviation and $R^{2}$)[28]; and radar vegetation index (RVI) values [30].
Pretrained Features: We leverage the Pretrained Remote Sensing Transformer (Presto) [31], a deep learning model pre-trained on diverse remote sensing time series data. We picked this remote sensing foundation model as it works at a pixel level as opposed to patch level, which would be difficult due to the irregular size of plots. We extract embeddings by passing ERA5 climate variables and Sentinel-1 timeseries data through the frozen Presto encoder, generating features that capture complex temporal patterns without requiring extensive labeled data. We utilize the embeddings from all timesteps (“temporal tokens”). A limitation of this model is that it can process at most 24 timesteps. We experiment with resampling the smoothed spline in the previous step with frequencies of 4, 7 and 10 days. This is done in order to increase the temporal resolution of the data where possible (when the total number of timesteps $\leq$ 24) or to accommodate longer temporal ranges (such as rows 11-12 in Table 2).
3.4 Classification Approach
We test two ensemble learning methods—Random Forest (RF) and Gradient Boosted Trees (GB)—for their proven effectiveness in agricultural classification tasks [32, 33]. These models effectively handle the heterogeneous feature sets while remaining computationally efficient. Following our dimensional framework, we structure three classification tasks: a) AWD vs. DSR vs. Control (null hypothesis: all practices on same dimension), b) DSR vs. PTR (sowing dimension), and c) AWD vs. CF (irrigation dimension). For each task, we split data into non-overlapping train/test sets (90:10 ratio) using stratified sampling. We optimize hyperparameters using Optuna [34] and evaluate performance using overall accuracy and F1 scores (macro and globally weighted averages). We compare against a baseline that randomly assigns classes based on class distribution proportions. This approach tests our hypothesis that separating water management practices along their natural dimensions (sowing and irrigation) produces more accurate results than attempting simultaneous multi-class classification (Fig. S1 for the complete workflow).
Classification Task Model Overall Accuracy F1 (weighted) F1 (macro)
AWD v/s DSR v/s PTR Baseline 33.47% (±3%) 33.38% (±3%) 33.27% (±3%)
RF 55.04% 54.39% 54.11%
GB 59.69% 59.62% 59.57%
DSR v/s PTR Baseline 56.8% (±3%) 56.3% (±3%) 50.01% (±3%)
RF* 75.19% 75.19% 71.76%
GB* 77.52% 77.82% 75.13%
AWD v/s CF Baseline 53.92% (±3%) 53.89% (±3%) 49.98% (±3%)
RF** 79.84% 78.76% 75.75%
GB** 80.62% 80.33% 78.07%
Table 1: Classification results for dimensional and combined approaches for the best performing models. RF = Random Forest, GB = Gradient Boosted Trees. Bold indicates best results per task. *=sampling frequency of 4 days, **=sampling frequency of 10 days
4 Results
4.1 Dimensional Classification Outperforms Combined Approach
Table 1 presents performance metrics across three classification tasks. The dimensional approach, which separates water management practices along sowing (DSR vs. PTR) and irrigation (AWD vs. CF) dimensions, significantly outperforms the combined three-class approach. When classified along their respective dimensions, both DSR and AWD achieve F1 scores of 78% and 80%, respectively. This result confirms our hypothesis that water management practices are more effectively detected when classified along their natural dimensions rather than being addressed simultaneously. However, as demonstrated in subsection 4.2, AWD classification relies predominantly on lag features, and the models fail to capture irrigation frequency directly, raising concerns about their generalizability. In contrast, DSR classification successfully exploits both lag and sowing features, indicating more robust pattern recognition.
Temporal Ranges Sowing Irrigation F1 (weighted, %)
Row Dates M J J A S O N D Lag (%) DSR PTR CF AWD ALL DSR AWD
1 May 1 - Aug 15 ~50% $\checkmark$ $\checkmark$ $>$ $\times$ 51.88 73.97* 73.48*
2 Jun 1 - Aug 30 ~50% $>$ $\checkmark$ $\checkmark$ $<$ 57.75 71.99 76.31
3 Jun 1 - Sep 15 ~50% $>$ $\checkmark$ $\checkmark$ $<$ 57.75 77.82* 76.31
4 Jun1 - Oct 15 ~60% $>$ $\checkmark$ $\checkmark$ $\checkmark$ 59.62 76.88 76.96
5 July 1 - Oct 15 ~50% $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 55.09 72.84 71.94
6 Aug 1 - Oct 15 ~33% $\times$ $\times$ $\checkmark$ $\checkmark$ 48.94 69.64 70.92
7 Aug 1 - Nov 15 ~50% $\times$ $\times$ $\checkmark$ $\checkmark$ 43.39 72.84 70.69
8 Sep 1 - Dec 15 ~50% $\times$ $\times$ $\checkmark$ $\checkmark$ 49.86 67.68 66.20
9 Oct 1 - Dec 15 ~33% $\times$ $\times$ $<$ $\checkmark$ 51.93 65.80 64.27
10 Aug 1 - Dec 15 ~60% $\times$ $\times$ $\checkmark$ $\checkmark$ 47.00 73.07 68.30
11 Jun 1 - Dec 15 ~33% $>$ $\checkmark$ $\checkmark$ $\checkmark$ 59.62 75.34** 79.12**
12 May 1 - Dec 15 ~90% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 55.29 73.46** 80.33**
Table 2: Effect of temporal range selection on accuracy. Features are generated for different temporal ranges to analyze the effect on outcomes by manipulating the availability of lag and practice features. Cell colors for months represent how much of the month is covered for that row (blank=none, green=full, orange=half). Values in bold indicate the top-2 performers for each classification task. Nominal sampling frequency of 7 days; *= a sampling frequency of 4 days, **=10days.
4.2 Temporal Features Drive Classification Performance
Our analysis reveals two distinct types of temporal features affecting classification performance:
Lag features capture differences in planting schedules between farms. PRANA collected an independent dataset documenting planting dates across practice types (Supp. Fig. S3), which reveals that while substantial variance exists in planting dates across the state, each practice type exhibits a considerably tighter distribution. These temporal offsets provide significant discriminatory power; however, this discriminatory capability depends on the persistence of such patterns over different seasons.
Practice features reflect water management patterns unique to each technique. Sowing features (active for DSR/PTR) are most distinct during the first weeks after planting, while AWD/CF irrigation patterns emerge later in the growing season, as a function of their irrigation frequency, which typically ranges from 4 to over 10 on average (Fig. 4).
To assess the relative importance of these features, we selectively constrained features to different temporal windows during model training (Table 2), thereby highlighting the differential prominence of each feature type. For sowing classification, models exposed to sufficient lag features (>~50%) and containing the majority of DSR and PTR sowing features (rows 1, 3, 10-11) demonstrate strong performance. When DSR features are absent (rows 4-9) or insufficient lag features are available (row 2), DSR classification performance deteriorates markedly. This pattern indicates that sowing models leverage both practice and lag features, conferring partial resilience to changes in cropping calendars. Furthermore, the optimal model appears robust to sampling frequency selection, as the top two performing models employ different sampling frequencies (f=4,7 for rows 3,4).
For irrigation classification, however, we observe strong dependence on lag features, with performance peaking when all features are available (row 11) and declining progressively as lag features are reduced (rows 11 >> row 8). Moreover, performance collapses in the absence of sowing features (rows 5-9), while models maintain reasonable performance even without irrigation features (rows 1-2). This indicates that AWD performance is predominantly driven by lag features and largely disregards irrigation features. This counterintuitive result suggests that AWD irrigation signals are not effectively captured with Sentinel-1A’s 12-day revisit frequency, which proves insufficient for detecting AWD irrigation cycles that typically occur every 4-10 days. This interpretation is further supported by evidence from rows 11-12, where increased sampling frequency to 10 days (i.e. lower temporal resolution incapable of capturing high frequency AWD events) yields optimal model performance.
4.3 Error Analysis Reveals Impact of Temporal Overlap
For both DSR and AWD classification, errors occur predominantly in samples with temporally similar planting dates (Fig. S4). This error pattern confirms that while our dimensional approach effectively addresses most planting date heterogeneity, classification remains challenging when practices exhibit similar temporal signatures. The elevated error rate in AWD classification in the presence of irrigation features [^3] These rows were additionally modeled with a higher sampling frequency of 4 days, yielding even poorer results. (Table 2 row 8-10) further corroborates our finding that current satellite revisit frequencies inadequately capture irrigation cycles.
4.4 Optimizing for Large-Scale Monitoring
For operational monitoring at scale, we tested various satellite data configurations. We found that ascending orbit data consistently outperforms descending orbit data for DSR classification, with no significant benefit from combining both orbits. Additionally, including climatology variables (temperature and precipitation) provides no performance boost (Fig. S5). These findings suggest that a streamlined approach using only ascending orbit data without climatology variables is optimal for large-scale monitoring, reducing computational requirements while maintaining classification performance.
4.5 Inference at scale
A key challenge in promoting sustainable agriculture is understanding adoption patterns of water-saving practices across large regions. The PRANA project trained 150,000 farmers on water-saving techniques, but tracking their implementation was impossible through traditional surveys. Our satellite-based approach enables monitoring at unprecedented scale, providing critical insights for food security policies. To apply our classification framework across Punjab, we first needed to delineate individual agricultural plots. With approximately 3.2 million hectares of rice cultivation and average plot sizes of 0.5-2 hectares, millions of individual fields required identification. We automated this process using the Fields of the World (FTW) algorithm [20], which processes bi-temporal Sentinel-2 imagery to identify field boundaries. Our workflow (Fig. S2) begins with acquiring cloud-free Sentinel-2 imagery from May and October 2024. After preprocessing and dividing the region into manageable tiles, we applied the FTW model to generate field polygons, filtering for size (2,000 $\text{m}^{2}$ to 10 hectares) to match typical smallholder farms. For each identified field, we processed Sentinel-1 time-series data and extracted embeddings from Presto to capture temporal signatures. To address the challenge of sparse training data (1,300 labeled plots within approximately 3 million agricultural plots extracted), we implemented an ensemble approach, selecting the top model from both RF and GB frameworks. These models’ predictions were aggregated by taking the modal value for each rice-growing plot, enhancing robustness and mitigating individual model biases, with the final results in Fig. 1.
Fig. 1 A. shows the spatial distribution of predictions of DSR practicing rice-fields for 2024. It is obvious from visual inspection that there is significant adoption heterogeneity across districts, and even within districts. We validated these results against Punjab government records, which tracked 253,328 acres of DSR cultivation in 2024 through their incentive program [35]. Our district-level predictions showed strong overall correlation with these records (Pearson=0.77, RBO= 0.77444Rank biased overlap (RBO) [36] with p=0.95, standard weighting), though with regional variations (Fig. 1 B). Some discrepancies were observed, with our model underestimating DSR adoption in Fazilka and Sri Muktsar Sahib while overestimating in Sangrur, Patiala, Barnala, Bathinda, and a few other districts, relative to government records. The high RBO indicates a good level of preservation of district rankings, particularly among the highest-ranked districts, complemented by a strong linear correlation between these estimates (Fig. 1 C). These differences may reflect both model limitations and potential inaccuracies in government data, as similar incentive programs in neighboring states have faced verification challenges [37]. For the majority of districts, the model overestimates DSR adoption, except in the top two districts, relative to the provided estimates. The higher RBO relative to Pearson correlation suggests that our model likely captures the relative ordering of districts than absolute values, making it valuable for prioritization exercises and policy applications where identifying relative standings across districts is paramount. Importantly, it enables targeting a sub-district or even at village or plot level. This large-scale application demonstrates how satellite monitoring can provide timely, comprehensive data on sustainable agriculture adoption, enabling more effective policy interventions to address water scarcity and food security challenges.
5 Limitations
Our approach, while effective, faces several constraints that merit consideration. First, the current Sentinel-1 revisit frequency of 12 days limits our ability to capture high-frequency irrigation variability characteristic of AWD practices. This temporal resolution challenge could be addressed by integrating data from ascending and descending orbits or incorporating forthcoming satellite missions such as Sentinel-1C, which would provide more frequent observations of irrigation cycles.
Second, field boundary extraction using the FTW algorithm is constrained by Sentinel-2’s spatial resolution, which imposes a minimum detectable plot size. This limitation is particularly significant in smallholder farming contexts where field sizes may fall below this detection threshold, potentially excluding smaller agricultural parcels from analysis.
Finally, our models’ performance relies heavily on lag features that capture temporal distinctions between different water management practices. The effectiveness of this approach depends on DSR practices maintaining sufficient temporal separation from other water management methods. Before generalizing these methods to other rice-growing regions globally or even temporally in the same region, additional validation studies must be conducted to determine whether these temporal patterns persist across diverse agroecological contexts and farming calendars.
6 Conclusion & Future Work
Our study demonstrates that satellite monitoring of rice water management practices is feasible, offering a powerful tool for tracking sustainable agriculture adoption. By classifying practices along separate sowing and irrigation dimensions, we achieved 78% F1-score in distinguishing direct seeded rice from traditional transplanted methods across highly variable planting dates.
This dimensional approach proved critical—DSR classification performed significantly better than attempting to classify all practices simultaneously. The framework successfully scaled to approximately 3 million plots across Punjab, with district-level predictions showing strong correlation (0.79) with government records. Our method requires no prior knowledge of planting dates, making it uniquely suited for real-world agricultural landscapes where cropping schedules vary widely.
While DSR detection performed robustly, monitoring irrigation practices like AWD remains challenging with current satellite revisit frequencies. The upcoming Sentinel-1C mission, expected in 2025, will provide the 6-day temporal resolution needed to better capture irrigation patterns.
This framework directly supports food security initiatives by enabling: 1) evidence-based resource allocation to promote sustainable water management, 2) measurement of intervention impacts at regional and farm scales, 3) targeted support for regions with low adoption rates, and 4) comprehensive monitoring without costly ground surveys
Future work should extend this approach to other major rice-producing regions facing water scarcity. By adapting the framework to diverse agricultural contexts, we can support sustainable intensification globally. Additional opportunities include developing region-specific models, integrating deep learning foundation models like Presto, and creating open data platforms for policymakers and researchers.
As climate change intensifies water scarcity challenges, scalable monitoring of sustainable water management will become increasingly vital for maintaining agricultural productivity and food security. This framework provides a foundation for tracking and promoting practices that allow farmers to produce more food with fewer resources while maintaining regional water security—directly supporting the Zero Hunger Sustainable Development Goal.
Appendix
Appendix A Data and Code Availability
The data and code that enables study replication shall be made available at the end of the peer-review period.
Appendix B PRANA Program
We focused our analysis on the Indian state of Punjab, both because of the scale of rice farming, and importance to national and global rice supply, as well as the existing program by The Nature Conservancy, the Promoting Regenerative and No-burn Agriculture (PRANA) project [^5] PRANA website, to promote water-saving rice practices as well and reduce crop residue burning. India is the world’s second biggest producer of rice [38], with the state of Punjab producing 10.42% of the country’s rice in 2024 at 144 million tons[11] representing roughly 3.7% of the world’s rice production capacity. Punjab’s prominence in rice production is heavily supported by irrigation infrastructure and high-yield farming practices but raises concerns about water sustainability due to the heavy reliance on groundwater systems. In Punjab, the adoption of the rice-wheat cropping system has led to a manifold increase in the irrigation water demand, and about 73% of it is met from groundwater [15], with the state experiencing a strong decline groundwater capacity[39] in the last few decades. A large part of this is due to the demand for water for rice cultivation, since the state of Punjab is in a semi-arid region with highly seasonal monsoon precipitation. Thus, unreliable surface water supplies coupled with excessive groundwater exploitation due to free electricity[13] has led to a long-term groundwater decline of 41.6 cm/yr in the state[12]. Han et al. [17] find that rice paddy planting regions increased significantly in northwestern India - one of the regions with the highest conversion from single- to double-cropping of rice.
The PRANA program conducted focused training sessions for ~150,000 farmers on a targeted scale to generate awareness, provide education, agronomic support and monitoring of farmers’ irrigation practices over the years 2023-2024, to enable adoption of both DSR and AWD methods. For both DSR and AWD, potential areas were identified based on the soil type and availability of required machinery. Field demonstrations was conducted in nearly every five villages to showcase the techniques locally. For AWD practice, 27,011 pipes were provided to the farmers pro bono and 26,997 hectares were covered under AWD using these pipes. Apart from these activities, paddy crop guides on agronomic practices were designed and provided to the farmers in local languages to build and strengthen adoption capacity. Fig. 2 shows field photos where these interventions were put into place in 2024.
B.1 Prediction in the Presence of Planting Date Heterogeneity
Previous studies that discriminated DSR from PTR relied on synchronizing SAR backscatter data with cropping calendars, enabling direct modeling of temporal signatures[10, 8, 9]. While effective, this approach is inherently limited in scalability, particularly in regions with unknown or heterogeneous planting dates.
Fig. A1: Planting date distributions across water management practices of interest.
As shown in Fig. A1, the planting dates in our dataset exhibit a roughly trimodal distribution: DSR plots are sown first, followed by control plots, and finally AWD plots. This temporal disparity spans up to 110 days between the first and last planting dates, introducing substantial temporal variability both between and within classes. This heterogeneous distribution naturally introduces temporal differences between classes, and where the distributions are wide enough - within classes. By not synchronizing cropping calenders across sample plots, we introduce lagged features—temporal offsets capturing the relative differences in planting dates—into our analysis. Lagged features enable the model to account for inter-class and intra-class variability in planting schedules, providing discriminatory power without requiring explicit cropping calendar data. Therefore lagged features can act as a proxy for cropping calenders as long as they exhibit low dispersion within a class. In practice, however, these features also reflect the limitations imposed by wide planting date distributions, which can span up to 2.5 months within individual classes like DSR and PTR as seen in Fig. A1.
Although planting and harvesting dates were available in our dataset, we intentionally excluded them from the modeling process to enhance the generalizability of our framework. This decision ensures that our approach remains applicable to regions where such information is unavailable. Additionally, farmer interviews revealed the presence of "third crops" (such as lentils, maize, or green manuring crops666a green manure is a crop specifically cultivated to be incorporated into the soil while still green) planted just before the kharif rice season. These crops introduce further variability in SAR backscatter signatures, complicating the detection of water management practices. Due to incomplete data on these practices, third-crop effects were not explicitly modeled but represent an additional source of temporal heterogeneity.
Fig. A2: Examples of hand-crafted features generated from SAR backscatter timeseries for different irrigation field types. Top-left: DSR and CF. Top-right: PTR and AWD. Bottom row: PTR with CF. Left subplots show the VV (blue) and VH (orange) timeseries for every pixel contained within a plot, and the dark blue and dark orange show the mean (with shaded regions denoting the standard deviation) VV and VH, respectively, per plot. Right subplots show the gaussian fit to the VV/VH curve and the radar vegetation index (RVI). The dotted lines in each left-subplot denote the planting date for that plot.
Appendix C Modeling Details
We experiment with ensemble-based learning methods including random forest (RF) and gradient-boosted trees (GB) to classify rice plots into one of 3 possible classes: AWD, DSR or PTR. Previous works have shown the suitability of these methods to LULC and agricultural practice mappings [32, 33, 40, 41]. Since we can reduce plot-level features to a point estimate, time-series modeling is possible with ensemble models like RF or GB. We use scikit-learn’s implementation for RF and LightGBM for GBs.
Hand-crafted Features: We generate a set of hand-crafted features from each S-1 band ($VV,VH,VV/VH$) per orbit, as the summary statistics for a rice-growing plot. For each band, we locate the positions of the first $K$ troughs and crests as well as the first $K$ inflection points. We extract the timestamp and amplitude at these points, measuring timestamp relative to the first date in the timeseries, in number of days. Experimentally, it was found that $K=3$ would cover all timeseries in the training dataset, with larger $K$s leading to missing values. We also featurize the total number of trough and crests, along with the mean, minimum and maximum amplitude values of each band per orbit. Following previous work [28], we fit a Gaussian kernel to the VV/VH ratio and employ its parameters (amplitude, day of peak, standard deviation and $R^{2}$ of kernel fit) as additional features. Additionally, we calculate the Radar Vegetation Index (RVI) as modified for S-1[30] as it has been shown to be useful in detecting crop height. The timeseries generation process for S-1 bands after processing and additional hand-crafted features are visualized in Figure A2.
Presto Features: In addition to hand-crafted features, we employ the Pretrained Remote Sensing Transformer (Presto) [31] as a temporal feature extractor. Presto is transformer-based deep learning model that has been pre-trained on remote sensing pixel-timeseries data, including Sentinel-1 SAR, Sentinel-2 multispectral, Dynamic Earth land use and land cover (LULC) labels, ERA5 temperature and precipitation datasets and, latitude & longitude. To extract embeddings, we freeze the Presto encoder, fix the Dynamic Earth class to ‘cropland’, and input both ERA5 variables and Sentinel-1 VV and VH bands from either ascending or descending orbits. Since we fit a spline to smoothed SAR timeseries, we can use arbitrary sampling frequencies to pseudo-assess the effect of sampling frequency. This is especially important as a limitation of the Presto encoder is that it cannot process more than 24 timesteps-in such cases we can increase the sampling duration to fit the entire timeseries. Therefore we choose $f$=4,7,10 days. Additionally, we extract embeddings from the encoder at each timestep  [^7] Presto takes a timestamp numbering for the month number, however we find that it gives better performance to use the week number instead of the month, since irrigation and sowing signal frequency is higher than a month, plus one feature for the location, generating a total of $1128\times(T\times 2)+$ 1888One set of embeddings for S1 and Dynamic World inputs embeddings. This process is encapsulated within the ‘Presto’ block of Fig. A6.
C.1 Training Pipeline
The full training pipeline is shown in Figure A6.
Fig. A3: Impact of orbit direction and climatology data (temperature and precipitation) on overall accuracy of DSR and AWD classification performance. Left: Bar plot grouped by orbit direction (ascending, descending, both) with confidence intervals highlighting variability within each task. Right: Bar plot showing the impact of climatology inclusion (yes, no) on OA, with confidence intervals highlighting variability within each task.
C.2 Inference Pipeline
The full inference pipeline is shown in Figure A7.
Appendix D Ablations
D.1 Effect of Sentinel-1 Orbit
For DSR classification, ascending orbits consistently outperformed descending orbits, as shown in Figure A3 (left). This suggests that radiometric properties of the descending orbit make it suboptimal for capturing DSR features in this region. The use of both orbits showed no significant advantage over ascending orbit alone, and the additional effort required to process both orbits may not be justified. For AWD classification, orbit direction had less impact, likely due to lag features driving performance regardless of orbit, while AWD-specific features were not captured effectively. This highlights that ascending orbits better capture DSR features compared to descending orbits in this study.
D.2 Effect of Climatology Data
High overlap in CIs for models with and without climatology data indicates no significant benefit from including these variables (Figure A3 , right). As the average performance without ERA5-Land data was slightly higher, and considering the effort required to acquire and integrate this dataset, there appears to be no inherent advantage to its inclusion for this region and task.
These findings suggest that consistent use of ascending orbits is optimal for DSR classification, while climatology data may be omitted without loss of performance, simplifying large-scale inference.
Appendix E Error Analysis
On the held out test set of 129 plots from our stratified split, we find that for DSR classification,
Given the distributed but nearly separable cropping calendar (Figure A1), it is useful to analyze the sources of classification errors with respect to the original classes. For the best-performing DSR classification model, Figure A4 (top-left) highlights the proportion of test set errors originating from each original class, showing a linear decrease in errors as planting dates increment from left to right. Similarly, Figure A4 (top-right) illustrates the proportion of incorrect samples and their original classes for AWD classification, with errors decreasing towards the left, for dates further from the AWD class.
For both tasks, errors predominantly stem from temporally similar samples (~33% of that class if incorrectly classified), suggesting that lag features may not provide sufficient discrimination in these cases and that practice features lack the necessary power. Confusion matrices (Fig. A4 bottom row) reveal that negative class scores dominate the results. For both classification tasks, the positive classes(DSR, AWD) achieves precision of 75% and recall of 67%.
Fig. A4: Error Analysis. Top-left: DSR classification error percentages on test set from original classes. Bottom-left: DSR classification confusion matrix from collapsed classes. Top-right: AWD classification error percentages on test set from original classes. Bottom-right: AWD classification confusion matrix from collapsed classes. Text on bars represents percentage of errors contributed from that class.
E.1 Feature Importance
Fig. A5: Feature importance: Using the gain attributed to each feature, the feature importance for LightGBM models are estimated for sowing (DSR) classification, left and irrigation (AWD) classification on the right.
For the best performing models for DSR and AWD classification, we estimate the feature importance as shown in Figure A5. For DSR classification, the best performing model was trained on the time period Jun 1- Sep 15 with a 4 day sampling frequency (Figure 4, row 3). Therefore the features that start with “time_*” refer to the time period in 4 day increments beginning on Jun 1999Since we pass in both S1 timesteps and a single fixed value for Dynamic World, the Presto Encoder generates: [0] lat-lon token, [1:T+1] temporal tokens, [T+1:2T+1] Dynamic World tokens. The top two features are therefore from the time period Jun 16-20, implying that the features in those regions were most useful in classification. From Figure A1, using the lens of sowing detection, we see that this refers to the time when the majority of PTR crops have been sowed (control/DSR has peaked and AWD has begun), which aligns with our conclusion that the sowing features are actually helpful.
For AWD classification the model was trained for the period May 1 to Dec 15 with a ten day sampling frequency (Figure 4, row 12). Similarly, the top two features (Figure A5 left) correspond to the time periods of June 20-30, and Sep 18-28. The former is roughly the same timeperiod as the top-2 features for DSR, while the latter likely captures some important lag feature, since all plots are at some stage of growth in September, regardless of when they were planted.
Appendix F Large-scale Inference
An important goal of this study is to understand the spatial extent of use of water-saving rice management practices in Punjab. Out of the 150,000 farmers who received training and support for the 2024 kharif season, it was infeasible to track the practices of all participants and their networks during the season to measure the efficacy of these interventions. The data collection efforts of PRANA were limited to only a small fraction of the potential market. Here we present our findings on scaling the model’s predictions across the state at a plot-level.
[…]
Thereafter, we follow the method laid out in subsection 4.5 with ascending orbit only, and use the best pretrained model for DSR classification, without any climatology inputs. The combination of pretrained models such as Presto and FTW, along with our dimensional training methodology that requires no cropping calendar information, allows us to apply this methodology to arbitrary spatial scales.
Fig. A6: Training and validation workflow used in the study. We utilize Copernicus Dataspace and GEE to acquire Sentinel-1 and ERA5 temporal data on a per plot basis, and to subsequently reduce them to plot level statistics. We then fit a spline, resample, generate hand-crafted features and embeddings from Presto over a given time period. The separability of water management classes is modeled using a Random Forest and Gradient Boosted Trees. The inset shows an example of the average S-1 VV and VH timeseries after smoothing and spline fitting, with confidence intervals at each point in time.
Fig. A7: Inference Workflow: Methodology followed during DSR classification predictions for 3 million polygons using S-1 timeseries. FTW[20] polygons are generated from bitemporal 4-bands S-2 imagery